H gas = 0c in tank A. H gas = 300 kelvin in tank B. Describe difference in kinetic energy

Grade:10

1 Answers

Arun
6 years ago
Kinetic energy depends on temp.
E = (3/2) * K * T           where T is in kelvin
Hence 
E{H(0°c)} = (3/2) * K* 273
E{H(27°c)} = (3/2) * K * 300